OpenStudy (anonymous):

Lesson 02.02: Properties of the Atom • Describe the properties and locations of protons, neutrons, and electrons in an atom. • Explain what the mass number and atomic number identify about an atom. • Use mass and atomic number to determine the number of each subatomic particle in an atom. Lesson 02.03: Isotopes and Weighted Averages • Identify the differences between isotopes of an element. • Calculate the weighted average mass of an element when given the percent abundance and mass of its isotopes.

OpenStudy (frostbite):

Fast question: you know all this right, we just recap it?

OpenStudy (anonymous):

yes i already know this! this is just review... I promise!!

OpenStudy (frostbite):

No it is more because if it is then I see no reason why you don't make the work and tell me about the properties and locations for Protons: Neutrons: Electrons: (might be a good way to remember it as you do an effort your self instead I just explain it)

OpenStudy (frostbite):

Then I check if they are correct.

OpenStudy (frostbite):

or something missing,

OpenStudy (anonymous):

Protons and neutrons are in the center of the atom, making up the nucleus. Electrons surround the nucleus.and have a negative charge. Protons have a positive charge. neutrons have a no net charge

OpenStudy (frostbite):

Perfect so far :)

OpenStudy (frostbite):

Think I go though the rest: Explain what the mass number and atomic number identify about an atom. Standard notation: \[\LARGE _{Z}^{A} \left[ element \right]\] Z = number of protons = number of electrons = atomic number A = nucleon number = Z + N N = number of neutrons. Identify the differences between isotopes of an element: Elements: Atom with a specific amount of protons. Isotopes: Elements with different nucleon number. (Z = constant) AKA N (the number of neutrons change). Calculate the weighted average mass of an element when given the percent abundance and mass of its isotopes. http://openstudy.com/study#/updates/4e94f79b0b8b7d552aecbb07